OpenStudy (narissa):

Select the proper equation in point-slope form for the given point and slope. Point: (5, 9); Slope: 2 (Points : 1) y - 9 = 2(x- 5) y - 5 = 2(x + 9) y + 9 = 2(x - 5) y - 9 = 2(x + 5)

pooja195 (pooja195):

\[\huge y-y_1=m(x-x_1)\] where m=slope (x1,y1)

OpenStudy (narissa):

yes i was watching a few videos on khan academy about this but i still am confused

pooja195 (pooja195):

On what?

OpenStudy (narissa):

khan academy its a website where they show you step by step math problems

pooja195 (pooja195):

Right lets do this step by step are you familiar with the point slope form ?

OpenStudy (narissa):

a little i tryed doing this by myself and i got b is that correct or am i way off

OpenStudy (narissa):

i mean A

pooja195 (pooja195):

thats correct